OpenStudy (anonymous):

2x^2-8/3x+3/2=2/3 solve equation

OpenStudy (anonymous):

@Arnab09

OpenStudy (anonymous):

is it like this? \[2x ^{2}-8/3 x +3/2 = 2/3\]

OpenStudy (anonymous):

yes but i figured it out. can you help me with different one then?

OpenStudy (anonymous):

ok, post the question

OpenStudy (anonymous):

{6x-5y=10 10x-5y=-10

OpenStudy (anonymous):

in this particular case, if you subtract the first equation from the second one, you will be left with 4x=-20, or x=-5 put that value of x in any of those 2 equations and you will get the value of y

OpenStudy (anonymous):

what about this one? (4x^2)^3 (-2x^5)(-7x^6) simplify

OpenStudy (anonymous):

@Arnab09

OpenStudy (anonymous):

you know the rule of indices?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

at the first part you are cubing x^2, so it will be x^6, cube -4 as well, it will be -64 x^6